1.Now is January 2016. Sophie will buy 50,000 pounds of lean hog in March 2016. To hedge the hog price, she buys a March 2016maturity lean hog contract at the future price of $1.20per pound. Suppose that when the contract matures in March 2016, the market price of hog turns out to be $1per pound. This future contract calls for purchase of 50,000 pounds

.a) Calculate Sophie's payoff from purchasing 50,000 pounds of hogs in the spot market in March 2016 (note: a payment for purchase is a negative payoff)

b) Calculate Sophie's payoff from the future market in March 2016;
c) Calculate Sophie's total payoff from both the spot and future markets in March 2016;

d) If the market price of hog goes up to $1.50 per pound in March 2016, what's Sophie's total payoff?

e) Draw the future contract payoff graph for Sophie.

Please do working it out step by step showing work, not through excel
Answer #1

a.

Spot Market price = $1 per pound

Qty to buy = 50000 pounds

Sophie's Payoff from Spotmarket = -50000*$1 = -$50000

b.

Future Market price = $1.20 per pound

Qty to buy = 50000 pounds

Market price = $1

Sophie's Payoff from Future market = 50000*($1-$1.2) = -$10000

c.

Total Payoff = -50000-10000 = -$60000

d.

Market Price =$1.5

Sophie's Payoff from Future market = 50000*($1.5-$1.2) = $15000

Sophie's Payoff from Spotmarket = -50000*$1.5 = -$75000

Total Payoff = -$75000+$15000 = -$60000